I have a word problem need help on.
In 2007, there were 13,300 student at college A, with a projected enrollment increase of 1000 students per year. In the same year, there were 26,800 students at college B, with a projected enrollment decline of 500 students per year. According to these projections, when will the colleges have the same enrollment? What will be the enrollment in each college at that time?
Thanks for your help.

Answer by dukesmart(20)

13,300+1000x = 26800 - 500x
1000x + 500x = 26800 - 13,300
1500x = 13500
x = 13500/1500
x = 9
year of 2016 and each college has 22,300 students.